Geary Manufacturing has assembled the data appearing in the next column pertaining to two products. Past experience has shown that the unavoidable fixed manufacturing overhead included in the cost per machine hour averages $10. Geary has a policy of filling all sales orders, even if it means purchasing units from outside suppliers. Total machine capacity is 50,000 hours.

Electric

Blender

Mixer

Direct materials

$ 6

$11

Direct labor

4

9

Manufacturing overhead

at $16 per hour

16

32

Cost if purchased from an

outside supplier

20

38

Annual demand (units)

20,000

28,000

Question

If 50,000 machine hours are available and Geary Manufacturing desires to follow an optimal strategy, it should

  • A.Produce 25,000 electric mixers and purchase all other units as needed.

  • B.Produce 20,000 blenders and 15,000 electric mixers, and purchase all other units as needed.

  • C.Produce 20,000 blenders and purchase all other units as needed.

  • D.Produce 28,000 electric mixers and purchase all other units as needed.
Answer #1

To find the correct answer, we need to calculate the necessary machine hours per unit.

Blender requires 1 hour per unit. (16/16)

Mixer requires 2 hours per unit. (32/16)

So, with 50,000 machine hours, it can either produce 50000 blenders or 25,000 mixers or a combination but some units will still need to be purchased.

Now, to separate fixed portion of OH cost which averages $10 per machine hour,

Variable portion of OH for Blender = (16 * 1) - (10 * 1) = $6 per unit

Variable portion of OH for Mixer = (16 * 2) - (10 * 2) = $12 per unit

Total VC for blender = 6 + 4 + 6 = $16 per unit

Total VC for mixer = 11 + 9 + 12 = $32 per unit

The purchase price from outside supplier of blender is $20. So, cost savings = $20 - $16 = $4. Cost savings per machine hour = $4 / 1 mach hr per unit = $4 per machine hour

The purchase price from outside supplier of mixer is $38. So, cost savings = $38 - $32 = $6. Cost savings per machine hour = $6 / 2 mach hr per unit = $3 per machine hour

As the cost savings of blenders is more, company should produce blenders. As the annual demand is 20,000 blenders, it should produce 20,000 blenders. It would consume 20,000 machine hours.

Remaining machine hours = 50,000 - 20,000 = 30,000 hours

Mixers that can be produced = 30,000 hrs/ 2 hrs per unit = 15,000 units.

So, correct answer = B.Produce 20,000 blenders and 15,000 electric mixers, and purchase all other units as needed.